# The de Jong fundamental group of a non-trivial abelian variety is non-abelian

### v2

@@ Abstract.
We show the de Jong fundamental group of any non-trivial abelian variety over a complete algebraically closed extension $C/\mathbb{Q}_{p}$ is non-abelian. Generalizing an argument for $\mathbb{P}^{1}_{C}$, we also show the de Jong fundamental groups of positive-dimensional quasi-projective rigid analytic varieties over $C$ (including abelian varieties) depend on $C$ and can be arbitrarily large.
@@ 1. Introduction
However, in the rigid analytic setting, there are more general notions of covering spaces that mix properties of topological and finite étale coverings. One such notion is that of de Jong coverings as introduced in [deJong.EtaleFundamentalGroupsOfNonArchimedeanAnalyticSpaces], and there is an associated de Jong fundamental group $\pi_{1,\mathrm{dJ}}(A^{\mathrm{rig}},0)$, a pro-discrete topological group whose continuous actions on discrete sets correspond to de Jong coverings of $A^{\mathrm{rig}}$. It is well known that de Jong fundamental groups can be much larger than classical étale fundamental groups: for example, although $\pi_{1,\mathrm{\acute{e}t}}(\mathbb{P}^{n,\mathrm{rig}}_{C})=\pi_{1,\mathrm{\acute{e}t}}(\mathbb{P}^{n}_{C})=\{1\}$, for $n\geq 1$ the de Jong fundamental group $\pi_{1,\mathrm{dJ}}(\mathbb{P}^{n,\mathrm{rig}}_{C})$ is large: a connected component over $C$ of the height $n+1$ Lubin-Tate tower furnishes, via the Gross-Hopkins period map, a tower of connected de Jong coverings of $\mathbb{P}^{n,\mathrm{rig}}_{C}$. For $y\in\mathbb{P}^{n,\mathrm{rig}}(C)$, this tower is classified by
@@
From a classical perspective, an abelian variety having a non-abelian fundamental group is surprising. On the other hand, it is perhaps not as surprising as the non-triviality of the fundamental group of $\mathbb{P}^{n,\mathrm{rig}}_{C}$ furnished already by the covers of Eq. 1, and we use these covers in our proof of Theorem 1. Note that, in [Howe.ThedeJongFundamentalGroupOfP1C], we also gave a short argument using geometric Sen theory to show that the fundamental group of $\mathbb{P}^{1,\mathrm{rig}}_{C}$ depends on the choice of $C$ and can be arbitrarily large. In §4, we generalize this result so that it applies, in particular, to all quasi-projective rigid analytic varieties (including abelian varieties).
For $n=\dim A$, we establish Theorem 1 by pulling back the cover of $\mathbb{P}^{n}_{C}$ in Eq. 1 along a generically étale map $A\rightarrow\mathbb{P}^{n}_{C}$ and then using geometric Sen theory to show the monodromy of this pullback remains non-abelian. The proof also implies that the monodromy is infinite, an observation we first learned from Sasha Petrov.
@@
This method of proof does not work for a general abeloid variety $A/C$ — indeed, such an $A$ may not admit any rational functions, in which case the covers of Eq. 1 cannot be used to produce any candidate non-abelian covers of $A$. It is thus natural to ask: is the de Jong fundamental group of a non-trivial abeloid non-abelian? We note that, for $A/C$ abeloid, the usual profinite étale fundamental group is “as expected,” i.e. equal to $T_{\hat{\mathbb{Z}}}A$ (see [BellovinCaiHowe.CharacterizingPerfectoidCoversOfAbelianVarieties, Appendix A]).
The key ingredients in the proof of Theorem 1 are the injectivity of the geometric Sen morphism for the Lubin-Tate tower (Lemma 2, which follows from a computation of the geometric Sen morphism given in [DospinescuRodriguezCamargo.JacquetLanglands]), and the observation that, if the monodromy is abelian, then the geometric Sen morphism is constant (Lemma 3). We treat these preliminaries in Section 2 and then use them to prove Theorem 1 in Section 3. In §4 we state and prove Theorem Theorem 2, the generalization of [Howe.ThedeJongFundamentalGroupOfP1C, Theorem 1] that was described in Remark 1.
@@
Our proof does not use the surjectivity of Eq. 1; all that is needed is the pointwise injectivity of the associated geometric Sen morphism (a “softer” fact than the actual monodromy computation for Eq. 1).
@@ Acknowledgements
We thank Rebecca Bellovin, Hanlin Cai, and Ben Heuer for helpful discussions about profinite étale covers of abelian varieties around [BellovinCaiHowe.CharacterizingPerfectoidCoversOfAbelianVarieties]. We thank Sasha Petrov for helpful comments on an earlier draft and for pointing out to us that the geometric Sen morphism can be used to show that, for $X/C$ a connected rigid variety, $x\in X(C)$, and $f:X\rightarrow\mathbb{P}^{n}_{C}$ a map that is not locally constant, the image of $\rho_{n}\circ f_{*}:\pi_{1,\mathrm{dJ}}(X,x)\rightarrow\mathrm{SL}_{n+1}(\mathbb{Q}_{p})$ must contain an infinite closed subgroup; this inspired our proof of Theorem 1. This work was supported in part by National Science Foundation grants DMS-2201112 and DMS-2501816.
@@ 2.2. Representations of the de Jong fundamental group
In particular, we can apply this to the $\rho_{n}$ of Eq. 1, whose associated torsor is a geometric connected component of the infinite level Lubin-Tate tower.
@@ 3. Proof of Theorem 1
4. Fundamental groups of quasi-projective rigid analytic varieties depend on $C$ and can be very big
@@
Suppose $X/C$ is a connected rigid analytic variety such that there exists a map $f:X\rightarrow\mathbb{P}^{n,\mathrm{rig}}_{C}$ and a smooth point $x\in X(C)$ with
$df_{x}:T_{X,x}\rightarrow T_{\mathbb{P}^{n,\mathrm{rig}}_{C},f(x)}$
surjective. Then, for $x_{0}$ in $X(C)$, if $S\subseteq\pi_{1,\mathrm{dJ}}(X,x_{0})$ is infinite and $2^{S}$ has cardinality less than that of $C$, then $S$ is not a set of topological generators for $\pi_{1,\mathrm{dJ}}(X,x_{0})$.
@@
  • Theorem 2 applies, e.g., if $X/C$ is connected reduced of dimension $n$ and admits a locally closed embedding into $\mathbb{P}^{m}$ (i.e. $X/C$ is quasi-projective). Indeed, in this case one can take the map $f$ to be the projection onto a generic $n$-plane $\mathbb{P}^{n}\subseteq\mathbb{P}^{m}$. In particular, it applies to abelian varieties.
@@
We take $f$ as in the statement of the theorem; by composing with an automorphism of $\mathbb{P}^{n}_{C}$, we may assume $f(x)=[0:0:..:1]=0\in\mathbb{A}^{n}(C)\subseteq\mathbb{P}^{n}(C)$. We may then choose a small ball $B$ in the smooth locus $X^{\mathrm{sm}}(C)$ containing $x$ such that $f|_{B}$ is a submersion onto its open image $f(B)\subseteq\mathbb{A}^{n}(C)$.
@@
Because $f|_{B}$ is a submersion, for $t\in f(B)$, the locus of points in $B$ where $\kappa_{f^{*}\mathbb{L}_{t}}$ vanishes identically is thus $f^{-1}(t)\cap B$. Since these loci are distinct for different $t\in f(B)$, we conclude that the local systems $f^{*}\mathbb{L}_{t}$ are non-isomorphic for all $t\in f(B)$. Since the cardinality of the open $f(B)\subseteq C^{m}$ is the same as the cardinality of $C$, we conclude by the same argument given in the proof of [Howe.ThedeJongFundamentalGroupOfP1C, Theorem 1].

### v3

@@ Abstract.
We show that the de Jong fundamental group of any non-trivial abelian variety over a complete algebraically closed extension $C/\mathbb{Q}_{p}$ is non-abelian. Generalizing an argument for $\mathbb{P}^{1}_{C}$, we also show that the de Jong fundamental group of any connected rigid analytic variety over $C$ admitting a non-constant map to $\mathbb{P}^{n}_{C}$ (e.g., an abelian variety) depends on $C$ and is big.
@@ 1. Introduction
However, in the rigid analytic setting, there are more general notions of covering spaces that mix properties of topological and finite étale coverings. One such notion is that of de Jong coverings as introduced in [deJong.EtaleFundamentalGroupsOfNonArchimedeanAnalyticSpaces], and there is an associated de Jong fundamental group $\pi_{1,\mathrm{dJ}}(A^{\mathrm{rig}},0)$, a pro-discrete topological group whose continuous actions on discrete sets correspond to de Jong coverings of $A^{\mathrm{rig}}$. It is well known that de Jong fundamental groups can be much larger than classical étale fundamental groups: for example, although $\pi_{1,\mathrm{\acute{e}t}}(\mathbb{P}^{n,\mathrm{rig}}_{C})=\pi_{1,\mathrm{\acute{e}t}}(\mathbb{P}^{n}_{C})=\{1\}$, for $n\geq 1$ the de Jong fundamental group $\pi_{1,\mathrm{dJ}}(\mathbb{P}^{n,\mathrm{rig}}_{C})$ is large: a connected component over $C$ of the height $n+1$ Lubin-Tate tower furnishes, via the Gross-Hopkins period map, a tower of connected de Jong coverings of $\mathbb{P}^{n,\mathrm{rig}}_{C}$. For $y\in\mathbb{P}^{n,\mathrm{rig}}(C)$, this tower is classified by a continuous homomorphism
@@
From a classical perspective, an abelian variety having a non-abelian fundamental group is surprising. On the other hand, it is perhaps not as surprising as the non-triviality of the fundamental group of $\mathbb{P}^{n,\mathrm{rig}}_{C}$ furnished already by the covers of Eq. 1, and we use these covers in our proof of Theorem 1. Note that, in [Howe.ThedeJongFundamentalGroupOfP1C], we used these covers for $n=1$ to give two constructions of many non-isomorphic local systems on $\mathbb{P}^{1,\mathrm{rig}}_{C}$, then deduced that the fundamental group of $\mathbb{P}^{1,\mathrm{rig}}_{C}$ depends on the choice of $C$ and can be big. In §4, we generalize both constructions to obtain the same result for the fundamental group of any connected rigid analytic variety over $C$ admitting a non-constant map to $\mathbb{P}^{n,\mathrm{rig}}_{C}$ (e.g., an abelian variety).
For $n=\dim A$, we establish Theorem 1 by pulling back the cover of $\mathbb{P}^{n,\mathrm{rig}}_{C}$ in Eq. 1 along a generically étale map $A\rightarrow\mathbb{P}^{n}_{C}$ then using geometric Sen theory to show the monodromy of this pullback remains non-abelian. The proof also implies that the monodromy is infinite, an observation we first learned from Sasha Petrov.
@@
This method of proof does not work for a very general abeloid variety $A/C$ — indeed, such an $A$ may not admit any rational functions, in which case the covers of Eq. 1 cannot be used to produce any candidate non-abelian covers of $A$. It is thus natural to ask: is the de Jong fundamental group of a non-trivial abeloid non-abelian? We note that, for $A/C$ abeloid, the usual profinite étale fundamental group is “as expected,” i.e. equal to $T_{\hat{\mathbb{Z}}}A$ (see [BellovinCaiHowe.CharacterizingPerfectoidCoversOfAbelianVarieties, Appendix A]).
The key ingredients in the proof of Theorem 1 are the injectivity of the geometric Sen morphism for the Lubin-Tate tower (Lemma 2, which follows from the computation of the geometric Sen morphism given in [DospinescuRodriguezCamargo.JacquetLanglands] or a more general computation in [Howe.GeometricSenAndKodairaSpencer]), and the observation that, if the monodromy is abelian, then the geometric Sen morphism must be constant (Lemma 3). We treat these preliminaries in Section 2 and then use them to prove Theorem 1 in Section 3. In §4 we state and prove Theorem 2, the generalization of [Howe.ThedeJongFundamentalGroupOfP1C, Theorem 1] that was described in Remark 1.
@@
Our proof of Theorem 1 does not use the surjectivity of Eq. 1; all that is needed is the pointwise injectivity of the associated geometric Sen morphism (a “softer” fact than the actual monodromy computation for Eq. 1).
@@ Acknowledgements
We thank Rebecca Bellovin, Hanlin Cai, and Ben Heuer for helpful discussions about profinite étale covers of abelian varieties around [BellovinCaiHowe.CharacterizingPerfectoidCoversOfAbelianVarieties]. We thank Sasha Petrov for helpful comments on an earlier draft and for pointing out to us that the geometric Sen morphism can be used to show that, for $X/C$ a connected rigid variety and $f:X\rightarrow\mathbb{P}^{n}_{C}$ a map that is not locally constant, the pullback of the Lubin-Tate local system along $f$ is non-trivial. This work was supported in part by National Science Foundation grants DMS-2201112 and DMS-2501816.
@@ 2.2. Representations of the de Jong fundamental group
In particular, we can apply this to the $\rho_{n}$ of Eq. 1, whose associated torsor is a geometric connected component of the infinite level Lubin-Tate tower. We note that, the associated $\mathbb{Q}_{p}$-local system $\mathbb{L}$, viewed as a sheaf on the $v$-site $\mathbb{P}^{n,\mathrm{rig}}_{C,v}$ (or on perfectoid objects in $\mathbb{P}^{n,\mathrm{rig}}_{C,\mathrm{pro\acute{e}t}}$), admits a construction by modifications of vector bundles on the Fargues-Fontaine curve as in [ScholzeWeinstein.BerkeleyLecturesOnPAdicGeometryAMS207, Lecture 24]. In particular, it sits in a short exact sequence involving the period sheaf $\mathbb{B}$ sending a perfectoid $S/\mathbb{P}^{n,\mathrm{rig}}_{C}$ to $\mathcal{O}(Y_{S^{\flat}})$ (for $Y_{S^{\flat}}$ as in [Fargues.GeometrizationOfTheLocalLanglandsCorrespondenceAnOverview, Definition II.1.15]). Explicitly, for $\varphi$ the Frobenius on $\mathbb{B}$ and $\theta:\mathbb{B}\rightarrow\mathcal{O}$ Fontaine’s map, we can write this short exact sequence as
(2) $0\rightarrow\mathbb{L}\rightarrow\mathbb{B}^{\varphi^{n+1}=p}\xrightarrow{b\mapsto\sum_{i=1}^{n+1}\theta(\varphi^{i-1}(b))\cdot z_{i}}\mathcal{O}(1)\rightarrow 0$
where here $\mathcal{O}(1)$ denotes the usual line bundle on $\mathbb{P}^{n,\mathrm{rig}}_{C}$ (viewed as a $v$-bundle) and $z_{1},\ldots,z_{n+1}$ are its global sections giving the usual homogeneous coordinates.
@@ 3. Proof of Theorem 1
4. Fundamental groups of quasi-projective rigid analytic varieties depend on $C$ and can be big
@@
Suppose $X/C$ is a connected rigid analytic variety such that there exists a non-constant map $f:X\rightarrow\mathbb{P}^{n,\mathrm{rig}}_{C}$ for some $n\geq 1$. Then, for $x_{0}$ in $X(C)$, if $S\subseteq\pi_{1,\mathrm{dJ}}(X,x_{0})$ is infinite and $2^{S}$ has cardinality less than that of $C$, then $S$ is not a set of topological generators for $\pi_{1,\mathrm{dJ}}(X,x_{0})$.
@@
  • Theorem 2 applies, e.g., if $X/C$ admits a locally closed embedding into $\mathbb{P}^{n}$ (i.e. $X/C$ is quasi-projective). In particular, it applies to abelian varieties.
@@
As in the proof of [Howe.ThedeJongFundamentalGroupOfP1C, Theorem 1], it suffices to produce a collection of non-isomorphic local systems indexed by a set of the same cardinality as $C$. We give two such constructions, generalizing the constructions of [Howe.ThedeJongFundamentalGroupOfP1C, Lemma 1] and [Howe.ThedeJongFundamentalGroupOfP1C, Lemma 2]. The first construction applies for any $C$ but requires a small restriction on $f$ or $X$; it is based on pulling back Lubin-Tate local systems along ramified maps and computing the geometric Sen morphism as in [Howe.ThedeJongFundamentalGroupOfP1C, Lemma 1]. The second construction applies whenever $C$ has cardinality greater than $2^{\mathbb{N}}$ (which suffices for the result since the statement of Theorem 2 is otherwise vacuous), and is based on computing extensions of the trivial local system by the Lubin-Tate local system and the infinite-dimensionality of the first cohomology of pullbacks of the Lubin-Tate local system (generalizing a very simple part of the argument in [Hansen.ACounterexample]).
First construction. Suppose given a non-constant map $f:X\rightarrow\mathbb{P}^{n}_{C}$, $n\geq 1$. This construction will apply under the further assumption that there exists a smooth point $x\in X(C)$ such that $f$ is a submersion at $x$, i.e
$df_{x}:T_{X,x}\twoheadrightarrow T_{\mathbb{P}^{n,\mathrm{rig}}_{C},f(x)}.$
Note that this is automatic if $n=1$ (since $df_{x}$ cannot be everywhere zero as the map is not locally constant). For general $n$, one can typically reduce from the existence of a general non-constant $f$ to the case where $f$ is a submersion at a smooth point $x$ as above by composing with projections: this reduction always works if $X$ is quasi-compact using [BhattHansen.TheSixFunctorsForZariskiConstructibleSheavesInRigidGeometry, Theorem 1] (which implies either the image does not contain all $C$-points, in which case one can project down to a hyperplane, or that the $df_{x}$ is generically surjective). More generally, if $X$ is of countable type, then the same argument works by applying [BhattHansen.TheSixFunctorsForZariskiConstructibleSheavesInRigidGeometry, Theorem 1] on a countable covering [^1] In general, the situation one needs to avoid to run this reduction argument is the case where $f$ is some type of space-filling map, i.e., a map that is nowhere a submersion but still surjective on $C$-points; such a construction seems plausible without the countability assumption. by quasi-compacts then invoking the Baire Category Theorem on $\mathbb{P}^{n}(C)$ to deduce that if it is not submersive at any $C$-point then it is not surjective on $C$-points.
Given $f$ and a point $x\in X(C)$ where $f$ is a submersion, by composing with an automorphism of $\mathbb{P}^{n}_{C}$, we may assume $f(x)=[0:0:..:1]=0\in\mathbb{A}^{n}(C)\subseteq\mathbb{P}^{n}(C)$. We may then choose a small ball $B$ in the smooth locus $X^{\mathrm{sm}}(C)$ containing $x$ such that $f|_{B}$ is a submersion onto its open image $f(B)\subseteq\mathbb{A}^{n}(C)$.
@@
Because $f|_{B}$ is a submersion, for $t\in f(B)$, the locus of points in $B$ where $\kappa_{f^{*}\mathbb{L}_{t}}$ vanishes identically is thus $f^{-1}(t)\cap B$. Since these loci are distinct for different $t\in f(B)$, we conclude that the local systems $f^{*}\mathbb{L}_{t}$ are non-isomorphic for all $t\in f(B)$. Since the cardinality of the open $f(B)\subseteq C^{m}$ is the same as the cardinality of $C$, we conclude.
Second construction As usual, the map $f$ is determined by the line bundle $\mathcal{L}=f^{*}\mathcal{O}(1)$ on $X$ and the $n+1$ global sections $s_{i}=f^{*}z_{i}$, $i=1,\ldots,n+1$ of $\mathcal{L}$ which generate $\mathcal{L}$. If the $s_{i}$ are not linearly independent, then the map factors through a hyperplane $\mathbb{P}^{n-1}_{C}\subseteq\mathbb{P}^{n}_{C}$; thus we may assume the $s_{i}$ are linearly independent.
Now, let us assume $C$ has cardinality greater than $2^{\mathbb{N}}$. Then, for $\mathbb{L}$ the rank $n+1$ dimensional Lubin-Tate local system on $\mathbb{P}^{n,\mathrm{rig}}_{C}$, we are going to show the set of isomorphism classes of rank $n+2$ local systems on $\mathbb{P}^{n,\mathrm{rig}}_{C}$ arising as extensions of $\mathbb{Q}_{p}$ by $\mathbb{L}$ has cardinality at least that of $C$. We claim it suffices to show that $\mathrm{Ext}^{1}(\underline{\mathbb{Q}_{p}},f^{*}\mathbb{L})$ has cardinality at least that of $C$. Indeed, in terms terms of representations of the de Jong fundamental group, the extension structures of a representation $V_{2}$ by a representation $V_{1}$ realized in a fixed representation $W$ are given by isomorphism classes of pairs consisting of an inclusion $V_{1}\rightarrow W$ and a map $W\rightarrow V_{2}$ inducing $W/V_{1}\xrightarrow{\sim}V_{2}$. Fixing bases, the cardinality of the collection of extensions structures on a fixed $W$ is thus bounded by that of $M_{\dim W\times\dim V_{1}}(\mathbb{Q}_{p})\times M_{\dim V_{2}\times\dim W}(\mathbb{Q}_{p})$, which is the same as that of $\mathbb{Q}_{p}$, which is $2^{\mathbb{N}}$. Thus, when $C$ has cardinality greater than $2^{\mathbb{N}}$, the Ext group can only have cardinality at least that of $C$ if the set of isomorphism classes of the local systems arising as extensions also has cardinality at least that of $C$.
To compute extensions, we note that we may work on the $v$-site $X_{v}$ (this is just a convenience; we could also work on $X_{\mathrm{pro\acute{e}t}}$ by working with a basis of perfectoid objects). Pulling back Eq. 2, $f^{*}\mathbb{L}$ sits in its defining short exact sequence on $X_{v}$,
$0\rightarrow f^{*}\mathbb{L}\rightarrow\mathbb{B}^{\varphi^{n+1}=p}\xrightarrow{b\mapsto\sum_{i=1}^{n+1}\theta(\varphi^{i-1}(b))\cdot s_{i}}f^{*}\mathcal{L}\rightarrow 0.$
In particular, for any $(c_{1},\ldots,c_{n+1})\in C^{n+1}$, we may pull back this extension along the map $\underline{\mathbb{Q}_{p}}\rightarrow f^{*}\mathcal{L}$ sending $1$ to $\sum c_{i}s_{i}$ to obtain an element of $\mathrm{Ext}^{1}(\underline{\mathbb{Q}_{p}},\mathbb{L})$. The global sections of $\mathbb{B}^{\varphi^{n+1}=p}$ on $C$ over the connected rigid analytic variety $X$ are given by $B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}$, where $B^{+}_{\mathrm{crys}}$ is the usual Fontaine period ring for $C$, thus this construction induces an embedding
$C^{n+1}/B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}\hookrightarrow\mathrm{Ext}^{1}(\underline{\mathbb{Q}_{p}},\mathbb{L}).$
Here we have used that the sections $s_{i}$ are linearly independent. The domain of this embedding can also be presented as the quotient of $C^{n}$ by an $n+1$-dimensional $\mathbb{Q}_{p}$-vector subspace — indeed, taking any codimension 1 subspace $C^{n}\subseteq C^{n+1}$, we have $B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}+C^{n}=C^{n+1}$ (e.g. by taking cohomology of the associated modification sequence on the Fargues-Fontaine curve, or by realizing the map from $B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}$ onto the quotient $C^{n+1}/C^{n}$ as the surjective logarithm map for a height $n+1$ one-dimensional formal group as in [ScholzeWeinstein.ModuliOfpDivisibleGroups]), so that
$C^{n+1}/B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}=C^{n}/B^{+,\varphi^{n+1}=p}_{\mathrm{crys}}\cap C^{n}$
where the intersection on the right is an $n+1$ dimensional $\mathbb{Q}_{p}$-vector space (e.g. by the taking cohomology of the same modification sequence or by realizing it as the rational Tate module of the same height $n+1$ one-dimensional formal group). In particular, this space has the same cardinality as $C$, so we conclude.
Remark 5.
It is known, e.g., for affinoid balls that the $\mathbb{F}_{p}$-cohomology is infinite dimensional in degree $1$ and depends on $C$. This can be used to show bigness of the fundamental group as in the second construction in the proof of Theorem 2. However, for a projective variety, the $\mathbb{F}_{p}$-cohomology is finite dimensional (it is the same as the algebraic étale cohomology) and the profinite étale fundamental group is the same as the algebraic one thus stable under change of $C$. Thus one needs $\mathbb{Q}_{p}$-local systems to get a result for projective varieties (including abelian varieties).